# Uncertainty-Based Extensible Codebook for Discrete Federated Learning in Heterogeneous Data Silos

### v1

@@ Abstract
Federated learning (FL), aimed at leveraging vast distributed datasets, confronts a crucial challenge: the heterogeneity of data across different silos. While previous studies have explored discrete representations to enhance model generalization across minor distributional shifts, these approaches often struggle to adapt to new data silos with significantly divergent distributions. In response, we have identified that models derived from FL exhibit markedly increased uncertainty when applied to data silos with unfamiliar distributions. Consequently, we propose an innovative yet straightforward iterative framework, termed Uncertainty-Based Extensible-Codebook Federated Learning (UEFL). This framework dynamically maps latent features to trainable discrete vectors, assesses the uncertainty, and specifically extends the discretization dictionary or codebook for silos exhibiting high uncertainty. Our approach aims to simultaneously enhance accuracy and reduce uncertainty by explicitly addressing the diversity of data distributions, all while maintaining minimal computational overhead in environments characterized by heterogeneous data silos. Through experiments conducted on five datasets, our method has demonstrated its superiority, achieving significant improvements in accuracy (by 3%–22.1%) and uncertainty reduction (by 38.83%–96.24%), thereby outperforming contemporary state-of-the-art methods. The source code is available at https://github.com/destiny301/uefl.
@@ 1 Introduction
Federated Learning (FL), well known for its capacity to harness data from diverse devices and locations—termed data silos—while ensuring privacy, has become increasingly crucial in the digital era. Despite its pivotal role in distributed computing, FL confronts a formidable challenge: the heterogeneity of data across different silos. Such diversity often results in a significant performance gap when integrating updates from local models into the global model. As illustrated in Figure 1, although local models can achieve rapid convergence and impressive performance on their own, the synthesized global model struggles to reach similar levels of performance. This issue is especially pronounced in FL due to its reliance on varied data sources.
@@
Recent studies (Ghosh et al., 2020; Agarwal et al., 2021; Liu et al., 2021; Kairouz et al., 2021; Zhang et al., 2022; Yuan et al., 2022) have made significant advancements in addressing data heterogeneity within Federated Learning (FL), with one notable approach being the use of discrete representations to enhance model robustness against minor data shifts. Nonetheless, this strategy struggles to generalize models to data silos exhibiting significant distributional differences. Furthermore, these methods face difficulties in adapting to unseen data distributions, as they typically require the entire model to be re-trained. Such constraints limit their flexibility in adapting to the dynamically changing data landscapes, posing challenges for their applicability in real-world scenarios.
Figure 1: Results on MNIST dataset after 5 rounds. In the case of heterogeneous data silos, not only does accuracy decline, but uncertainty also sees an increase of approximately 50%.
@@
Moreover, we identify a critical issue impacting the model’s performance across diverse data silos: uncertainty. As depicted in Figure 1, the global model’s accuracy not only deteriorates, but its uncertainty also trends upwards, signaling increased prediction instability, even in the face of potentially favorable accuracy metrics. To address these challenges, we introduce Uncertainty-Based Extensible-codebook Federated Learning (UEFL), a novel methodology that explicitly distinguishes between data distributions to improve both accuracy and uncertainty.
@@
Specifically, our design features an advanced codebook comprising a predetermined number of latent vectors (i.e. codewords), and employs a discretizer to assign encoded image features to their closest codewords. These codewords, serving as latent features, are then transmitted to subsequent layers for processing. The codewords undergo dynamic training to adjust and align with the latent features produced by the image encoder. Notably, our codebook is extensible, facilitating the easy addition of new codewords. Initially, codewords are uniformly distributed across all data silos; however, additional codewords may be selectively allocated to manage data from particular distributions, thereby ensuring explicit differentiation between them. Given the privacy constraints inherent in FL that limit direct data access, we incorporate an uncertainty evaluator based on Monte Carlo Dropout. This tool identifies data from distinct distributions characterized by high uncertainty. Throughout the training phase, our UEFL method systematically distinguishes data from varied distributions and initiates new codewords to expand the codebook, continuing until all data distributions are adequately represented. In the initial training cycle, only shared codewords are utilized. Subsequent cycles take advantage of the fully trained image encoder, allowing for the nearly perfect initialization of new codewords and facilitating rapid adaptation to new distributions. As a result, our UEFL model can accommodate data from even previously unseen distributions with fewer communication rounds, demonstrating its potential applicability in enhancing other FL algorithms.
@@
  • We identify a significant increase in model uncertainty across silos with diverse data distributions within the federated learning (FL) context, highlighting the challenge of data heterogeneity.
  • To address this heterogeneity, we introduce an extensible codebook approach that distinguishes between data distributions by stepwise mapping them to distinct, trainable latent vectors (i.e. codewords). This methodology allows for efficient initialization of newly added codewords using a K-means algorithm, closely aligning with the ground truth distributions and enabling rapid convergence during codebook training.
  • We propose a novel data-driven FL framework, named Uncertainty-Based Extensible-codebook Federated Learning (UEFL), which merges the extensible codebook with an uncertainty evaluator. This framework iteratively identifies data from diverse distributions by assessing uncertainty without requiring direct data access. It then processes this data by initializing new codewords to complement the existing codebook, ensuring that each iteration focuses on training the expandable codebook, which rapidly converges, thus allowing UEFL to adapt seamlessly to new data distributions.
  • Our empirical evaluation across five datasets demonstrates that our approach significantly reduces uncertainty by 38.83%-96.24% and enhances model accuracy by 3%-22.1%, evidencing the effectiveness of UEFL in managing data heterogeneity in FL.
@@ 2.1 Federated Learning
Federated learning (Konečnỳ et al., 2016; Geyer et al., 2017; Chen et al., 2018; Hard et al., 2018; Yang et al., 2019; Ghosh et al., 2020) represents a cutting-edge distributed learning paradigm, specifically designed to exploit data and computational resources across edge devices. The Federated Averaging (FedAvg) algorithm (McMahan et al., 2017), introduced to address the challenges of unbalanced and non-IID data, optimizes the trade-off between computation and communication costs by reducing the necessary communication rounds for training deep networks. The field of Federated Learning (FL) grapples with numerous statistical challenges, central among them the heterogeneity of data. Several methodologies (Zhao et al., 2018; Li et al., 2018, 2019; Kalra et al., 2023) have been developed to address this pivotal issue. PMFL (Zhang et al., 2022) approaches the heterogeneity challenge by drawing inspiration from meta-learning and continual learning, opting to integrate losses from local models over the aggregation of gradients or parameters. DisTrans (Yuan et al., 2022) enhances FL performance through train and test-time distributional transformations, coupled with a novel double-input-channel model architecture. Meanwhile, FCCL (Huang et al., 2022) employs knowledge distillation during local updates to facilitate the sharing of inter and intra domain insights without compromising privacy, and utilizes unlabeled public data to foster a generalizable representation amidst domain shifts. Additionally, contributions from various domains, such as the discrete approach to addressing heterogeneity by Liu et al. (Liu et al., 2021), provide further inspiration and valuable perspectives for our research endeavors.
@@ 3.1 Overall Architecture
Figure 2: UEFL flowchart. In the first iteration, all latent are mapped to initialized shared codewords by the discretizer $\theta_{D}$. In the next iterations, UEFL identifies data from heterogeneous distributions with the uncertainty evaluator, and complements new codewords to enhance the codebook. Uncertain data can select not only newly added codewords but also shared codewords.
Figure 2 illustrates the workflow of our UEFL. Consider multiple data distributions $\mathcal{D}_{1},\mathcal{D}_{2},...,\mathcal{D}_{M}$, with data samples $x\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ denote the input image’s height, width, and channel count, respectively, drawn from these $M$ distributions. Upon distributing the global model to local clients, data samples undergo local encoding via a shared encoder $\theta_{E}$ into feature representations $z\in\mathbb{R}^{h\times w\times c}$, with $h$, $w$, and $c$ representing the features’ shape. Subsequently, these features are reshaped into vectors $z\in\mathbb{R}^{l\times c}$, where $l$ is the number of tokens, and divided into $s$ segments $z_{i}\in\mathbb{R}^{l\times\frac{c}{s}},\forall i$, with $s$ indicating the segment count. Each segment is mapped to the closest codeword in the codebook via a discretizer $\theta_{D}$, then reassembled into complete vectors for classification. The classifier $\theta_{C}$ then deduces the class for the input data, completing the forward processing sequence as follows,
@@
Upon each iteration’s conclusion, assessing uncertainty via Monte Carlo Dropout becomes crucial, given the inaccessibility of data in FL due to privacy concerns. By evaluating uncertainty against a pre-established threshold, we identify data from heterogeneous distributions. When such data are detected, we augment the codebook with new codewords, initializing them using centroids of the generated image features, as delineated in Algorithm 1. This process leverages the fully trained encoder from prior iterations, employing K-means to ensure the new codewords are well-aligned with the actual data distribution, thus facilitating swift convergence in training. Consequently, this approach necessitates minimal communication rounds for the next iterations.
3.2 Extensible Codebook
To effectively manage heterogeneous data, we design an extensible codebook, beginning with a minimal set of codewords and progressively enlarging this set through a superior initialization strategy that benefits from our UEFL framework. This strategy facilitates stepwise mapping of diverse data distributions to distinct codewords. Starting with a larger codebook can introduce uncertainty in codeword selection due to the concurrent training of multiple codewords.
Similar to VQ-VAE (Van Den Oord et al., 2017), we employ latent vectors as codewords, initially setting up a compact shared codebook $c^{(0)}\in\mathbb{R}^{n\times\frac{c}{s}}$, where $n$ denotes the codebook size, initialized with a Gaussian distribution for utilization across all data silos. To associate the $i$th feature vector $z_{i}$ with a corresponding codeword $c_{i}$, the discretizer calculates the distance to all codewords, subsequently selecting the nearest one as follows.
$c_{i}=\operatorname*{arg\,min}_{j\in{1,2,...,n}}||z_{i}-c_{j}||_{2}$ (4)
Following each iteration’s uncertainty assessment, we identify which silos necessitate additional codewords for enhanced prediction accuracy. Suppose in the subsequent $t$th iteration, we need another $n$ codewords $c^{(t)}\in\mathbb{R}^{n\times\frac{c}{s}}$, then we initialize them via K-means rather than Gaussian distribution since the trained encoder after the first iteration can generate superior latent features. To optimize codebook usage, data silos demonstrating lower performance in the prior iteration are enabled to select codewords from both the newly incorporated $c^{(t)}$ and the initial $c^{(0)}$.
K-means Initialization. After the first iteration, the encoder, now trained, generates image features that closely mirror the ground truth distribution. Initializing codewords with the centroids of these features ensures alignment with the ground truth, expediting codebook training. Figure 3 illustrates this concept: gray points represent features from the trained encoder, clustered according to their data distributions. While direct data access is restricted, differentiation by uncertainty allows us to identify and utilize the centroids of these clusters via K-means for codeword initialization. This strategy hugely reduces the number of training rounds required for model convergence in subsequent iterations. Moreover, we refresh the shared codebook at each iteration.
@@ 3.2 Extensible Codebook
        Broadcast $\theta^{t}$ to all clients
@@
           Initialize accessible codewords
@@
           Map latent features to discrete codewords: $c=\operatorname*{arg\,min}_{j\in{1,2,...,n}}||z-c_{j}||_{2}$
           Predict with coded vectors: $p=f_{\theta_{E}}(c)$
           Compute codewords loss: $\mathcal{L}_{code}=||sg(c)-z||^{2}_{2}+\beta||c-sg(z)||^{2}_{2}$
@@
     Evaluate uncertainty: $e_{k}=\sum_{c}p\log p$
@@
Segmented Codebooks. For complex datasets, a finite set of discrete codewords might not fully capture the diversity of image features. To bolster the robustness of our methodology, we dissect features into smaller segments—using factors like 2 or 4—to pair them with multiple codewords, thus covering the entirety of a feature vector as illustrated in Figure 3. This segmentation exponentially increases the codeword pool to $n^{2}$ or $n^{4}$, ensuring a robust representation capacity without necessitating a large-scale increase and permitting efficient K-means-based initialization. This design minimizes runtime overhead associated with larger codebooks.

### v4

@@ Abstract
Federated learning (FL), aimed at leveraging vast distributed datasets, confronts a crucial challenge: the heterogeneity of data across different silos. While previous studies have explored discrete representations to enhance model generalization across minor distributional shifts, these approaches often struggle to adapt to new data silos with significantly divergent distributions. In response, we have identified that models derived from FL exhibit markedly increased uncertainty when applied to data silos with unfamiliar distributions. Consequently, we propose an innovative yet straightforward iterative framework, termed Uncertainty-Based Extensible-Codebook Federated Learning (UEFL). This framework dynamically maps latent features to trainable discrete vectors, assesses the uncertainty, and specifically extends the discretization dictionary or codebook for silos exhibiting high uncertainty. Our approach aims to simultaneously enhance accuracy and reduce uncertainty by explicitly addressing the diversity of data distributions, all while maintaining minimal computational overhead in environments characterized by heterogeneous data silos. Extensive experiments across multiple datasets demonstrate that UEFL outperforms state-of-the-art methods, achieving significant improvements in accuracy (by 3%–22.1%) and uncertainty reduction (by 38.83%–96.24%). The source code is available at https://github.com/destiny301/uefl.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: In the case of heterogeneous data silos, the global model of regular FL performs poorly compared to local models. By discretizing different domains into distinct latent spaces, our UEFL improves both accuracy and uncertainty. The reported values in the figure represent the average accuracy and uncertainty across the various data silos.
@@
Federated Learning (FL), well known for its capacity to harness data from diverse devices and locations—termed data silos—while ensuring privacy, has become increasingly crucial in the digital era, particularly with the explosion of data from mobile sources. Despite its pivotal role in distributed computing, FL confronts a formidable challenge: the heterogeneity of data across different silos. Such diversity often results in a significant performance gap when integrating updates from local models into the global model. In Figure 1, we compare the mean accuracy of local FL models with that of the global model after integration when addressing data silos with different distributions. While local models may perform impressively within their own data domains, the aggregated global model often struggles to achieve similar performance levels after synthesizing updates from these varied data sources. This issue is especially pronounced in FL due to its reliance on varied data sources.
@@
Recent studies (Ghosh et al., 2020; Agarwal et al., 2021; Liu et al., 2021; Kairouz et al., 2021a; Zhang et al., 2022; Yuan et al., 2022) have made significant advancements in addressing data heterogeneity within FL, with one notable approach being the use of discrete representations to enhance model robustness against minor data shifts. Nonetheless, this strategy struggles to generalize models to data silos exhibiting significant distributional differences. Furthermore, these methods face difficulties in adapting to unseen data distributions, as they typically require the entire model to be re-trained. Such constraints limit their flexibility in adapting to the dynamically changing data landscapes, posing challenges for their applicability in real-world scenarios.
@@
Moreover, we identify another critical issue impacting the model’s performance across diverse data silos: increased uncertainty, as shown in Figure 1. The global model’s accuracy not only deteriorates, but its uncertainty also trends upwards, signaling increased prediction instability. To address these challenges, we introduce Uncertainty-Based Extensible-codebook Federated Learning (UEFL), a novel methodology that explicitly distinguishes between data distributions to improve both accuracy and uncertainty.
@@
Specifically, our design features an advanced codebook comprising a predetermined number of latent vectors (i.e. codewords), and employs a discretizer to assign encoded image features to their closest codewords. These codewords, acting as latent representations, are passed to subsequent layers for processing. The codewords are dynamically trained to align with the latent features generated by the image encoder. To mitigate performance degradation when integrating local models from data silos with varying distributions, we initialize a small, shared codebook for all clients. Additional specific codewords are then introduced for individual client use, ensuring explicit differentiation between them. Since the initial codebook is small and requires only a few extensions, the final size remains compact, minimizing the associated computational overhead. Given the privacy constraints in FL, which restrict direct data access, we incorporate an uncertainty evaluator using Monte Carlo Dropout. This evaluator identifies data from diverse distributions, marked by high uncertainty. During training, UEFL systematically distinguishes between these varied distributions and dynamically adds new codewords to the codebook until all distributions are sufficiently represented. In the initial training cycle, shared codewords are randomly initialized. However, in subsequent cycles, the fully trained image encoder is leveraged to initialize new codewords using K-means, aligning them more closely with the data distribution and facilitating faster adaptation to various distributions. As a result, our UEFL model can accommodate data from previously unseen distributions with fewer communication rounds, making it applicable for enhancing other FL algorithms.
@@
  • We identify a significant increase in model uncertainty across silos with diverse data distributions in FL, highlighting the challenge of data heterogeneity.
  • To address this heterogeneity, we introduce an extensible codebook approach that distinguishes between data distributions by stepwise mapping them to distinct, trainable latent vectors (i.e. codewords). This methodology allows for efficient initialization of newly added codewords using a K-means algorithm, closely aligning with the training data feature distributions and enabling rapid convergence during codebook training.
  • We propose a novel data-driven FL framework, named Uncertainty-Based Extensible-codebook Federated Learning (UEFL), which merges the extensible codebook with an uncertainty evaluator. This framework iteratively identifies data from diverse distributions by assessing uncertainty without requiring direct data access. It then processes this data by initializing new codewords to complement the existing codebook, ensuring that each iteration focuses on training the expandable codebook, thus allowing UEFL to adapt seamlessly to new data distributions.
  • Our empirical evaluation across various datasets demonstrates that our approach significantly reduces uncertainty by 38.83%-96.24% and enhances model accuracy by 3%-22.1%, evidencing the effectiveness of UEFL in managing data heterogeneity in FL.
@@ 2.1 Federated Learning
Federated learning (Konečnỳ et al., 2016; Geyer et al., 2017; Chen et al., 2018; Hard et al., 2018; Yang et al., 2019; Ghosh et al., 2020; Ji & Luo, 2025; Luo & Ji, 2025) represents a cutting-edge distributed learning paradigm, specifically designed to exploit data and computational resources across edge devices. The Federated Averaging (FedAvg) algorithm (McMahan et al., 2017), introduced to address the challenges of unbalanced and non-IID data, optimizes the trade-off between computation and communication costs by reducing the necessary communication rounds for training deep networks. FL faces numerous statistical challenges, with data heterogeneity being one of the most critical. In real-world applications, data collected across different clients often varies significantly in terms of distribution, feature space, and sample sizes.
Several methodologies (Zhao et al., 2018; Li et al., 2018, 2019; Kalra et al., 2023) have been developed to address this pivotal issue. PMFL (Zhang et al., 2022) approaches the heterogeneity challenge by drawing inspiration from meta-learning and continual learning, opting to integrate losses from local models over the aggregation of gradients or parameters. DisTrans (Yuan et al., 2022) enhances FL performance through train and test-time distributional transformations, coupled with a novel double-input-channel model architecture. Meanwhile, FCCL (Huang et al., 2022) employs knowledge distillation during local updates to facilitate the sharing of inter and intra domain insights without compromising privacy, and utilizes unlabeled public data to foster a generalizable representation amidst domain shifts. Additionally, the discrete approach to addressing heterogeneity by (Liu et al., 2021), provide further inspiration and valuable perspectives for our research endeavors.
@@ 3.1 Overall Architecture
Figure 2: UEFL flowchart. In the first iteration, all latent are mapped to initialized shared codewords by the discretizer $\theta_{D}$. In the next iterations, UEFL identifies data from heterogeneous distributions with the uncertainty evaluator, and complements new codewords with K-means initialization to enhance the codebook. Clients with high uncertainty can select not only newly added codewords but also shared codewords.
Figure 2 illustrates the workflow of our UEFL. Consider multiple data distributions $\mathcal{D}_{1},\mathcal{D}_{2},...,\mathcal{D}_{M}$, with data samples $x\in\mathbb{R}^{H\times W\times D}$, where $H$, $W$, and $D$ denote the input image’s height, width, and channel count, respectively, drawn from these $M$ distributions. Upon distributing the global model to local clients, data samples undergo local encoding via a shared encoder $\theta_{E}$ into feature representations $z\in\mathbb{R}^{h\times w\times d}$, with $h$, $w$, and $d$ representing the features’ shape. Subsequently, these features are reshaped into vectors $z\in\mathbb{R}^{l\times d}$, where $l$ is the number of tokens, and divided into $s$ segments $z_{i}\in\mathbb{R}^{l\times\frac{d}{s}},\forall i$, with $s$ indicating the segment count. Each segment is mapped to the closest codeword in the codebook via a discretizer $\theta_{D}$, then reassembled into complete vectors for classification. The classifier $\theta_{C}$ then deduces the class for the input data, completing the forward processing sequence as follows,
@@
At the end of each iteration, assessing uncertainty through Monte Carlo Dropout is essential, given the privacy constraints of FL, which limit direct access to client data. By evaluating uncertainty against a pre-established threshold, we identify data from heterogeneous distributions. When such data are detected, we augment the codebook with $v$ new codewords initialized by K-means. These newly generated codewords are then exclusively accessible to the corresponding heterogeneous clients, updating the codebook size for the $k$th client from $v_{k}$ to $v_{k}+v$, as described in Algorithm 1. This process leverages the fully adapted encoder from previous iterations, utilizing K-means to ensure the new codewords are closely aligned with the actual data distribution, thereby facilitating faster convergence during training. Additionally, since the extended codewords are specific to individual client data and are not included in the integration with other local models, our method ensures that latent features from different distributions remain explicitly differentiated. Consequently, the global model performs better after integration, effectively handling data heterogeneity.
@@
  Randomly initialize model parameters $\theta$ and codebook with $v$ codewords
  Initially assign uncertainty for all clients to be zero: $e_{k}=0,\forall k$
@@
        Broadcast $\theta$ to all clients
@@
           if $e_{k}>(1+\gamma)\min_{\forall j\in 1,2,...,K}(e_{j})$ then
              K-means initialize another $v$ codewords and add them to codebook
              Update accesible codewords size for clients with high uncertainty: $v_{k}\leftarrow v_{k}+v$
           end if
@@
           Discretize latent features to codewords $c_{i}$, where $i=argmin_{j\in{1,2,...,v_{k}}}||z-c_{j}||_{2}$
           Predict with coded vectors: $p=f_{\theta_{C}}(c_{i})$
           Compute codewords loss: $\mathcal{L}_{code}=||sg(c_{i})-z||^{2}_{2}+\beta||c_{i}-sg(z)||^{2}_{2}$
@@
     Evaluate uncertainty for each client with integrated model: $e_{k}=\sum p\log p$
@@
3.2 Extensible Codebook
While discretization mitigates data heterogeneity (see Appendix A for theoretical analysis), it struggles with distributions exhibiting significant variations. To effectively handle such heterogeneous data, we design an extensible codebook, beginning with a minimal set of codewords and progressively enlarging this set through a superior initialization strategy that benefits from our UEFL framework. This strategy facilitates stepwise mapping of diverse data distributions to distinct codewords. Starting with a larger codebook can introduce uncertainty in codeword selection due to the concurrent training of multiple codewords.
Similar to VQ-VAE (Van Den Oord et al., 2017), we employ latent vectors as codewords, initializing a compact shared codebook with $v$ codewords $c\in\mathbb{R}^{v\times\frac{d}{s}}$, where $v$ represents the size of the initial codebook. The codewords are initialized using a Gaussian distribution and shared across all data silos. After each iteration’s uncertainty assessment, we determine which silos require additional codewords to improve prediction accuracy, and we extend the codebook accordingly for these silos by adding $v$ more codewords. The newly added codewords are initialized using K-means, leveraging the encoder’s improved latent features from the prior iteration to better align with the underlying data distribution. To optimize codebook usage, data silos that demonstrated lower performance in the previous iteration are allowed to select codewords from both the newly added codewords and the original shared codebook. Typically, the codebook only requires 1-3 extensions until all clients reach low uncertainty levels. The server updates the codebook by computing the average of codewords across the clients that utilize those specific codewords.
For a given iteration, if the codebook size for the $k$th client is $v_{k}$, the feature vector $z$ is associated with a codeword $c_{i}$ by the discretizer, which computes the distance between $z$ and all available codewords, selecting the nearest one as follows,
$i=\operatorname*{arg\,min}_{j\in{1,2,...,v_{k}}}||z-c_{j}||_{2}$ (4)
K-means Initialization. After the first iteration, the adapted encoder produces image features that more accurately reflect the distribution of the training data. Instead of relying on random initialization methods like Gaussian distribution, we initialize new codewords using the centroids of these features, obtained through K-means clustering. This approach expedites codebook training by providing a more informed starting point for the new codewords, allowing them to better align with the underlying data structure. As a result, this initialization strategy facilitates faster convergence and improves the model’s ability to adapt to varying data distributions across silos. This strategy hugely reduces the number of training rounds required for model convergence (Details in Appendix J).
Segmented Codebooks. For complex datasets, a finite set of discrete codewords might not fully capture the diversity of image features. To bolster the robustness of our methodology, we dissect features into smaller segments to pair them with multiple codewords, thus covering the entirety of a feature vector. This segmentation exponentially increases the codeword pool, ensuring a robust representation capacity without necessitating a large-scale increase and permitting efficient K-means-based initialization. This design minimizes runtime overhead associated with larger codebooks.